In a simple random sample of 150 households, the mean number of computers was 2.13. Assume the population standard deviation is σ = 0.45. Construct a 95% confidence interval for the mean number of computers.
Answer #1

Solution :

Given that,

Sample size = n = 150

Z/2 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (0.45 / 150)

Margin of error = E = 0.07

At 95% confidence interval estimate of the population mean is,

- E < < + E

2.13 - 0.07 < < 2.13 + 0.07

2.06 < < 2.20

(2.06 , 2.20)